OpenStudy (anonymous):

log x + log(x − 48) = 2 solve for x

OpenStudy (anonymous):

log x(x-48) = 2 x(x-48) = 10^2 (assuming log to base 10) x^2 - 48x - 100 = 0 solve for x

OpenStudy (anonymous):

can you continue?